# Sensor-fusion based Prognostics Framework for Complex Engineering Systems Exhibiting Multiple Failure Modes

### v2

@@ Abstract
Complex engineering systems are often subject to multiple failure modes. Developing a remaining useful life (RUL) prediction model that does not consider the failure mode causing degradation is likely to result in inaccurate predictions. However, distinguishing between causes of failure without manually inspecting the system is nontrivial. This challenge is increased when the causes of historically observed failures are unknown. Sensors, which are useful for monitoring the state-of-health of systems, can also be used for distinguishing between multiple failure modes as the presence of multiple failure modes results in discriminatory behavior of the sensor signals. When systems are equipped with multiple sensors, some sensors may exhibit behavior correlated with degradation, while other sensors do not. Furthermore, which sensors exhibit this behavior may differ for each failure mode. In this paper, we present a simultaneous clustering and sensor selection approach for unlabeled training datasets of systems exhibiting multiple failure modes. The cluster assignments and the selected sensors are then utilized in real-time to first diagnose the active failure mode and then to predict the system RUL. We validate the methodology using a simulated dataset of systems exhibiting two failure modes and on NASA turbofan degradation dataset.
@@
Multiple Failure Modes , Remaining Useful Life , Sensor Fusion , Functional Principal Component Analysis , Mixture of Gaussian Regression , Adaptive Sparse Group Lasso , Expectation Maximization
journal: Reliability Engineering & System Safety
@@
{highlights}
We propose a two-stage prognostic framework for multi-sensor systems with multiple failure modes.
Offline: An unsupervised method labels failure events and ranks sensor importance.
Online: Sensor features are fused for failure mode diagnosis, then key sensors predict RUL.
Our method achieves $80\%$ clustering accuracy and $<3\%$ RUL prediction error in simulations.
In NASA turbofan data, early predictions outperform benchmarks while reducing input by $30\%$.
@@ 1 Introduction
Prognostic models aim to estimate the remaining useful life (RUL) of critical components and systems. They are important because they drive condition-based maintenance decisions that aim to increase equipment availability, reduce downtime, and prevent unexpected failures. Generally, prognostic models can be classified into two categories: (1) data-driven models that focus on modeling sensor data and assume that the data is directly (or indirectly) correlated with the underlying physical degradation process, and (2) physics-based models that rely on first principles and require a thorough understanding of the underlying physics driving the degradation process. While physics-based approaches are known for their high accuracy, data-driven approaches are growing in popularity, especially for modeling complex systems. Several reasons are driving this trend. First, complex systems often exhibit complex physics of failure that are difficult to characterize using first principles. The second reason is the rise of Machine Learning and Artificial Intelligence models, which when trained properly can be very powerful prognostic tools. However, many data-driven prognostics models assume that complex systems exhibit a single predominant failure mode that can be used to accurately predict the system’s RUL. In practice, complex engineering systems often exhibit multiple failure modes and thus require more sophisticated modeling approaches to account for the different ways/modes a system can fail.
In this paper, we focus on developing a prognostic methodology for complex systems monitored by multiple sensors that exhibit multiple failure modes. We design an integrated framework comprised of (1) a sensor selection algorithm, (2) a dimensionality reduction and data fusion methodology, and (3) a RUL prediction model. Our methodology relies on a historical dataset to simultaneously diagnose observed failure events and select the subset of the most informative sensors associated with each failure mode. One of the unique aspects of our methodology relates to the sensor selection model, which uses an unsupervised approach to identify the dominant failure modes. Unlike most of the existing works, we assume that the labels of the failure events are unknown a priori. This assumption introduces some unique challenges around model estimation. These challenges are not trivial, especially when they are considered in conjunction with sensor selection.
@@
Our methodology is comprised of two parts, an offline component and an online prognostic model. In the offline part, we develop a novel algorithm based on the Expectation-Maximization (EM) algorithm that simultaneously labels the failures in the historical training set and selects the optimal sensors for each failure mode. In the online part, we utilize real-time data to perform online diagnosis of the predominant failure mode and RUL prediction. During the online learning phase, Multivariate Functional Principle Component Analysis (MFPCA) is utilized for feature extraction and dimensionality reduction. We then use the nearest neighbor classifier to diagnose the active failure mode in the real-time signals before fitting a weighted functional regression model using the extracted features. The main contributions of this paper are threefold:
  • First, we propose a feature extraction methodology comprised of fusing signals from multiple sensors originating from a set of systems exhibiting multiple failure modes.
  • Second, we propose an optimization methodology for simultaneous failure mode labeling and sensor selection for any given unlabeled historical dataset. We call this as the “offline sensor selection” step.
  • Finally, we provide a two-step sequential approach to leverage this knowledge of failure labels and informative sensors to (1) perform active failure mode diagnosis and (2) subsequent RUL prediction on systems degrading in real-time. The paper mentions this as the “online diagnosis and prognostics” step.
@@
The remainder of the paper is outlined as follows. In Section 2, we review the relevant literature associated with this research area. In Section III, we outline our proposed multi-failure mode prognostics framework. In Section 4, we present our offline sensor selection scheme that utilizes an EM algorithm. In Section 5, we present our online diagnostics and prognostics method. It discusses a strategy to fuse informative sensors for real-time diagnosis of failure modes and subsequent prediction of RUL. Our methodology is demonstrated through two case studies – one using a simulated dataset in section 6 and the other NASA turbofan engine dataset in section 7. We conclude by summarizing our framework and key findings in Section 8.
@@ 2 Literature Review
Data-driven prognostic models utilize a variety of approaches. A significant number of these rely on Artificial Intelligence tools, such as neural networks to predict component lifetime [1, 2]. Some approaches have focused on directly modeling degradation-based signals that are derived from raw sensor data. Examples include random coefficient models [3], relevance vector machines [4], stochastic processes such as the Brownian motion process [5], the Gamma process [6], and hidden Markov models [7], and similarity-based prognostic models [8]. Most of these models have been developed for modeling a single sensor or degradation-based signal. Extending these approaches to applications involving multiple sensors is not straightforward.
Fusion algorithms have been leveraged to model multivariate sensor data. According to [9], fusion is usually performed at the decision level or the data level. Decision-based fusion involves combining the results of multiple prediction models in an ensemble fashion [10]. Conversely, data fusion methodologies focus on aggregating data from multiple sensors to develop a single (univariate) health index, which can then be modeled using the conventional methods discussed above. Most of the works in this area are characterized by their data fusion approach. Examples include principal component analysis (PCA) [11, 12, 13], maximum likelihood [14], deep learning [15], and state space models [16, 17, 18]. Most data fusion methodologies tend to aggregate all the sensor data. In practice, some sensors may not be correlated with the underlying failure process. Thus, integrating signals from these sensors into a data fusion algorithm introduces noise and may impact RUL prediction accuracy. To address this issue, several works have incorporated sensor selection algorithms. For example, [14] and [9] employed adaptive lasso to penalize non-informative sensors. [13] combined location-scale regression with group regularization technique to perform sensor selection. While these techniques can select an optimal subset of sensors for monitoring, they still assume that system failure is caused by a single predominant failure mode.
Most complex engineering systems can fail due to multiple failure modes. Competing risk models have been a popular approach for modeling multiple failure modes. Traditionally, competing risk models rely on statistical approaches. For example, [19] uses the Cox proportional hazard model (PHM) [20] with the Weibull baseline hazard rate to develop a mixture of Weibull PHM. For an extensive review of statistical competing risk models, we refer the reader to [21]. Deep learning has also been used to develop competing risk models. These models include convolutional neural network (CNN) [22], recurrent neural networks [23], and LSTM [24]. Recently, [25] developed a deep learning methodology to model complex survival problems, such as recurrent hospitalizations due to kidney failure or pancreas graft. For more examples, we refer the reader to [26], which provides a review of machine learning methodologies for survival analysis. Competing risk models often output survival/hazard rates that are sometimes difficult to interpret in engineering applications. Consequently, many researchers have adopted methods that explicitly model sensor data to predict RUL.
The literature that focuses explicitly on RUL prediction of systems with multiple failure models can be loosely divided into two groups. The first group uses deep learning approaches while the second group utilizes more statistical methods. Deep learning models such as [27] and [28] develop feature extraction methodologies based on convolutional neural networks (CNN) for modeling systems with multiple failure modes. [29] developed a Gaussian model whose mean is a Bayesian deep neural network that maps sensor data to the RUL. In [30], the authors used a support vector machine to diagnose the active failure mode using features extracted from a deep belief network. Following diagnosis, a particle filter is employed to predict RUL. [31] trained an LSTM-based RUL prediction model for all present failure modes. The choice of prediction model is based on a deep CNN-based classifier trained with physics-informed failure labels. While deep learning approaches have been successful at predicting RUL, their parameters are often difficult to interpret. Additionally, they require significantly larger datasets for training compared to statistical methods. In [32], the authors derived a composite health index for each failure mode and used a random coefficients-based polynomial degradation model to predict the system’s RUL. In their model training/estimation, the authors assumed that the labels of the failure modes were given. In fact, this assumption has been widely used in the literature and is often supported by visual inspection of the data [27, 30, 32] or available domain knowledge [31]. [33] assumes that a subset of the data is labeled and utilizes a similarity-based approach for classifying the rest of the unlabeled failure modes in the training set.
@@
Our methodology differs from previous works in that we consider a scenario where the failure labels are completely unknown. Furthermore, we assume that for each failure mode, there exists a (not necessarily disjoint) subset of informative sensors that are useful for prognostics. Our objective is to design an unsupervised prognostic methodology that synthesizes signals from multiple sensors in a systematic fashion. We assume the availability of a historical repository of system failures whose degradation was monitored from the moment an incipient fault occurred up to system failure using multiple sensors. The methodology assumes that all systems fail independently of each other. In addition, we assume that only one type of failure mode can occur in the entire lifetime of the system. While the failure labels are unknown, the total number of potential failure modes for the system is assumed to be known.
@@ 3 Overview of the Prognostics Framework
This section describes our multiple failure mode prognostics framework, which is summarized in the flowchart in Figure 1. This framework is comprised of two parts:
Offline Sensor Selection: The first part is performed offline and primarily focuses on utilizing a training set of historically observed degradation signals to perform sensor selection. Sensor selection is important because not all sensors are informative for prognostics, and the inclusion of non-informative sensors in a prognostics algorithm can compromise the accuracy of RUL predictions. When performing sensor selection, it is computationally convenient to represent each sensor by a parsimonious set of features that capture the system dynamics. Therefore we further divide this step into the following substeps:
@@
    Feature Extraction: It involves a variant of the traditional Functional Principal Component Analysis (FPCA) framework to account for multiple failure modes in the unlabeled data.
@@
    Optimization: This involves formulating an optimization model to simultaneously label the failure modes and select the informative sensors for each failure mode.
Feature Extraction: There are many ways to extract parsimonious features from complex and multivariate data. In our framework, we choose to use a Functional Principal Component Analysis (FPCA) framework due to its flexibility and extensibility to various types of functional data, allowing us to efficiently capture the underlying patterns and variability in complex, high-dimensional datasets. FPCA models the variability of temporal data observations around a mean function. However, the presence of multiple failure modes can cause heterogeneity in the magnitude and/or sign of the signals’ rate of change, compromising the effectiveness of FPCA as a prognostic tool. Therefore, the first step of our offline sensor selection algorithm is to extract features using a Covariate-Adjusted FPCA (CA-FPCA) approach [34].
CA-FPCA is an extension of conventional FPCA that allows modeling functional data while accounting for the effects of explanatory variables, known as covariates. In our context, the covariates are used to represent the failure modes. One of the unique contributions of our approach is the assumption that the failure modes are not known a priori. However, CA-FPCA assumes that covariates are observable. We note that the way each sensor responds to different failure modes can be highly variable. For example, a sensor may be more sensitive to one failure mode than another resulting in heterogeneity in the rate of change of its corresponding sensor signals. Furthermore, a sensor signal may increase in magnitude in response to one failure mode, and decrease for another. To address the lack of observable covariates, we first perform regular FPCA on the signals from each sensor. Then, we perform K-means clustering on the FPC scores for each sensor individually, using the resulting cluster labels as the covariate values to perform CA-FPCA.
Optimization: After feature extraction the sensor signals are transformed into parsimonious CA-FPC-scores, which serve as predictors for a Mixture of Gaussian (MGR) model used to model the natural logarithm of the time-to-failure, $\ln{TTF}$, as a function of the CA-FPC-scores. This model is motivated by the assumption that the functional relationship between the $\ln{TTF}$ and the CA-FPC scores is dependent on the latent failure mode. We fit the parameters of the MGR model using the Expectation-Maximization (EM) algorithm. In the E-step, we obtain a soft clustering of the ((((CA-FPC-scores, $\ln{TTF})$ tuples into their respective failure modes. In the M-step, we incorporate the Adaptive Sparse Group Lasso (ASGL) penalty into the regression model to remove groups of features associated with non-informative sensors. From now on, we refer to this model as MGR-ASGL. The output of the offline sensor sensor selection step of our methodology are (1) failure mode labels for each sample in the training set and (2) a subset of informative sensors for each failure mode.
Online diagnosis and Prognostics: The second part of our methodology consists of an online step used to process and analyze data from units operating in the field. The online step consists of diagnosing the active failure mode of the fielded unit and then predicting the RUL of the fielded unit. In the diagnostic step, features are extracted from the real-time data generated by all selected sensors using Multivariate FPCA (MFPCA), which is an extension of FPCA to multivariate functional data. We use MFPCA to generate a parsimonious set of fused signal features from multivariate sensor signals; hereafter referred to as MFPC-scores. Then, we classify the active failure mode of the fielded unit by assigning to the unit the failure mode most frequently represented by its K-nearest neighbors. Once this step is complete, we recalculate the MFPC scores using only the informative sensors associated with the diagnosed failure mode of the unit. In the prognostic step, we fit a functional regression model using the MFPC scores that predict the RUL of the unit. In the following sections, we formalize our prognostics framework.
Figure 1: Flowchart of the multiple failure mode prognostics framework
4 Offline Sensor Selection Algorithm

### v3

# Sensor-fusion based Prognostics for Deep-space Habitats Exhibiting Multiple Unlabeled Failure Modes

@@ Abstract
Deep-space habitats are complex systems that must operate autonomously over extended durations without ground-based maintenance. These systems are vulnerable to multiple, often unknown, failure modes that affect different subsystems and sensors in mode-specific ways. Developing accurate remaining useful life (RUL) prognostics is challenging, especially when failure labels are unavailable and sensor relevance varies by failure mode. In this paper, we propose an unsupervised prognostics framework that jointly identifies latent failure modes and selects informative sensors using only unlabeled training data. The methodology consists of two phases. In the offline phase, we model system failure times using a mixture of Gaussian regressions and apply a novel Expectation-Maximization algorithm to cluster degradation trajectories and select mode-specific sensors. In the online phase, we extract low-dimensional features from the selected sensors to diagnose the active failure mode and predict RUL using a weighted regression model. We demonstrate the effectiveness of our approach on a simulated dataset that reflects deep-space telemetry characteristics and on a real-world engine degradation dataset, showing improved accuracy and interpretability over existing methods.
@@
Deep Space Habitats , Multiple Failure Modes , Remaining Useful Life , Sensor Fusion , Failure Diagnosis , Prognostics
journal: Acta Astronautica
@@ 1 Introduction
A deep-space habitat is a complex engineering system (CES), composed of multiple interacting subsystems that must work together to support human life and mission operations over long durations. These subsystems such as the environmental control and life support system, energy and power systems, thermal control, robotic agents, and structural components etc., must function reliably in harsh and isolated environments, where failure of any component can lead to mission-critical consequences [1, 2, 3, 4]. The complexity of such systems arises not only from the tight integration of these subsystems but also from their heterogeneous origins. Components are often developed by different manufacturers and tested independently, making it difficult to anticipate how they will behave once integrated. Additionally, the deep-space environment introduces operating conditions such as microgravity, radiation, and thermal extremes that differ significantly from terrestrial test settings. These novel conditions can give rise to failure modes that have never been observed during isolated ground-based validation. As a result, building a comprehensive library of failure modes becomes essential for predicting degradation behaviors (i.e., prognostics) during the mission. However, due to limited diagnostic access and the lack of labeled failure data, this library must often be inferred from unlabeled system observations.
In addition to the challenge of missing labels, deep-space habitats face another difficulty that limits their prognostic capability. These systems are typically equipped with a large number of sensors that generate in situ telemetry data to continuously monitor system health. However, analyzing this data presents several issues. Not all sensors are informative for predicting remaining useful life (RUL), and many may be redundant or affected by noise. Furthermore, in systems with multiple failure modes, sensor informativeness is often mode-specific. A sensor that is strongly predictive under one failure mode may be uninformative under another (see Figure 1). This variability complicates the task of selecting a reliable subset of sensors for prognostic modeling, especially with unknown failure modes.
In this paper, we propose a data-driven prognostics framework for deep-space systems that are monitored by many sensors and subject to multiple, unknown failure modes. Our methodology consists of two key stages: (1) an offline sensor selection step and (2) an online diagnosis and prognostics step. These stages align with two mission stages: during early deployment, the habitat operates under limited supervision to collect historical sensor data for initializing the model; after this setup phase, autonomous health monitoring and prediction begin.
Figure 1: Example of a deep space habitat with multiple subsytems. Health of each subsystem is monitored by a subset of onboard sensors. The subset of informative sensors vary for different failure modes.
@@
Main Contributions. A unique aspect of our approach is that we do not assume failure mode labels are known in the historical training data (degradation signals). This reflects realistic space mission constraints, where diagnosing past failures requires expert involvement and post-mission analysis. To address this, we model failure times using a mixture of Gaussian regressions, where regressors are features extracted from sensor data. We develop a novel Expectation-Maximization (EM) algorithm that simultaneously labels failure modes and selects the most informative sensors for each mode. In the online phase, we use real-time sensor data to first diagnose the dominant failure mode before predicting RUL. We apply Multivariate Functional Principal Component Analysis (MFPCA) to extract low-dimensional, yet informative features from the mode-specific sensors identified in the offline phase. We then classify the active failure mode using a nearest-neighbor approach and apply a weighted functional regression model to predict RUL. The key technical contributions of this paper can be summarized as:
  • 1.
    We develop a feature extraction methodology that fuses multivariate sensor data from systems exhibiting multiple failure modes into compact and informative representations.
  • 2.
    We propose a failure-mode-aware sensor selection approach using an Expectation-Maximization algorithm that jointly clusters unlabeled failure events and selects informative sensors for each failure mode.
  • 3.
    We present an integrated online framework that uses real-time data to (1) diagnose the active failure mode and (2) predict RUL using a mode-specific regression model.
  • 4.
    We demonstrate the effectiveness and robustness of the proposed methodology through two case studies relevant to deep-space habitats: (1) a controlled simulation designed to reflect key telemetry challenges, and (2) a real-world NASA turbofan engine degradation dataset that mimics sensor-driven subsystem degradation in long-duration missions.
@@
The remainder of the paper is organized as follows. Section 2 reviews related work on RUL prediction and multi-failure mode diagnostics. Section 3 introduces our proposed prognostics framework. Section 4 describes the offline sensor selection method. Section 5 details the online diagnosis and RUL prediction approach. Section 6 and Section 7 present two case studies—a simulated dataset and the NASA turbofan engine dataset, respectively. We conclude the paper in Section 8 with a discussion of future work.
@@ 2 Literature Review
Approaches for RUL prediction of CES are typically categorized as either physics-based or data-driven. Physics-based approaches [5, 6] utilize first principles to mathematically model the underlying physics of system degradation. For example, [5] used a two-stage nonlinear differential equation to model the rate of air pressure changes inside a space habitat. They use an Extended Kalman Filter (EKF) to estimate the parameters of this model and predict RUL by using the same EKF to forecast when the habitat pressure becomes untenable. In general, physics-based approaches typically develop state-space models of system degradation and use a Bayesian filter such as the Kalman Filter or the Particle Filter for real-time estimation and prediction of model parameters. Several examples of this are highlighted by [7]. The advantages of physics-based approaches include high prediction accuracy and limited need for data. However, their requirement for thorough knowledge of the physics behind system failure renders them difficult to develop, given the growing system complexity.
Conversely, data-driven approaches attempt to map sensor data to the system RUL using statistical models, machine learning, or hybrid approaches. Since sensors are monitoring the system over a prolonged period, they generate continuous, high-volume data called “streaming data.” The approaches developed for single-sensor streaming data include both statistical methods that model univariate degradation signals using exponential models [8, 9], random-coefficient regression [10], Brownian motion [11], Gamma process [12], or hidden Markov model [13] and machine learning methods that use deep learning [14, 15] or relevance vector machine [16]. However, single-sensor approaches are not scalable to the scenario where multiple sensors monitor systems. Under this scenario, sensors produce multivariate streaming data. For this type of data, authors have developed sensor fusion algorithms that operate at either the decision level or the data level [17]. Decision-level fusion consists of systematically combining RUL estimates from different models, whereas data-level fusion synthesizes the multivariate sensor data into a univariate health indicator, which can be modeled using standard univariate techniques. Existing data fusion approaches include Principal Component Analysis [18, 19, 20], maximum likelihood methods [21], extreme learning machine [22], deep learning [23, 24], state-space models [25, 26, 27], and logistic regression [28].
While effective, many data fusion methods rely on aggregating all available sensor signals. In practice, not all sensors are informative for a given failure mechanism, and fusing irrelevant signals can introduce noise and reduce prediction accuracy. To address this, some studies incorporate sensor selection. For instance, adaptive lasso has been used to eliminate uninformative sensors [17, 21], and group regularization techniques have been applied for joint selection and modeling [20]. However, these approaches typically assume a single dominant failure mode.
In reality, most CES can fail through multiple failure modes, reflecting the structure and function of their subsystems. Competing risk models have been widely used to represent such systems. Traditional approaches often rely on statistical models, such as the Cox proportional hazards model with Weibull baselines [29, 30]; a comprehensive review is provided in [31]. More recently, deep learning has also been used to model competing risks [32]. For example, both [33] and [34] both employ deep recurrent neural networks for recurrent event survival analysis with competing risks.
Competing risk models often produce survival or hazard rates that can be difficult to interpret in engineering contexts. As a result, many researchers have shifted toward methods that directly model sensor data to predict RUL, though few focus on multiple failure modes. It is common for researchers to develop general RUL prediction methodologies and then apply them to datasets with multiple failure modes. For example, [35] proposed a correlation and relative entropy feature engineering framework for complex systems and [36] proposed a deep convolutional neural network. In both cases, the researchers validated their methodologies on datasets with multiple failure modes. However, they do not consider multiple failure modes in an explicit manner. Failure to do so could result in highly biased predictions. Therefore, the development of RUL prediction methodologies that account for multiple failure modes is growing. [37] used a support vector machine for failure mode diagnosis with features extracted from a deep belief network, followed by RUL prediction using a particle filter. [38] trained distinct LSTM models for each failure mode. Once a physics-informed CNN classifier diagnoses the active failure mode, the appropriate LSTM model is utilized. Furthermore, [39] fit random coefficient models to failure mode-specific health indices. However, these articles do not incorporate a sensor selection algorithm and they rely on labeling of the failure modes determined either through visual inspection of domain knowledge. Consequently, researchers have begun addressing these gaps. [40] adopted a semi-supervised graph-based approach for feature extraction from partially labeled failures and used elastic net functional regression for sensor selection and RUL prediction. More recently, [41] proposed a time-series clustering algorithm to identify latent failure modes, followed by LSTM-based RUL prediction. Furthermore, [42] combined mixture of (log)-location scale regression with deep learning to predict RUL in a scenario with unknown failure modes. However, in both papers, sensor selection was limited to excluding sensors with obviously non-informative signals, such as constant or low-variance signals.
@@
Our methodology differs from prior work by addressing a setting in which failure mode labels are entirely unknown. We assume that each failure mode is associated with a (not necessarily disjoint) subset of sensors that are informative for RUL prediction. The goal is to develop an unsupervised prognostic framework that systematically fuses multivariate sensor signals for RUL estimation. We assume access to a historical repository of system run-to-failure trajectories, where degradation was monitored from the onset of an incipient fault to final failure using multiple sensors. Furthermore, we assume that each system fails independently and experiences only one failure mode during its lifetime. Finally, while the failure labels are unobserved, we assume that the total number of potential failure modes is known. In the following sections, we detail our prognostics framework.
Table 1: Comparison of Prognostic Methodologies in the Literature
Prognostic Sensor Sensor Multiple Unlabeled Model Methodology Fusion Selection Failure Modes Failure Modes Interpretability [5, 6, 7] N/A [8, 9, 10, 11, 12, 13] [18, 19, 20, 21, 23] Partial [21, 20, 43] [39, 44, 37, 38] Partial [40] Partial Partial [41, 42] Partial Partial [32, 45, 33, 34, 46] [44, 35, 47, 36] Ours
@@ 3 Overview of the Prognostics Framework
Figure 2: Overview of the proposed prognostics framework for deep-space habitats
This section describes our prognostics framework, developed for autonomous health monitoring in deep space habitats. A flowchart of the framework is shown in Figure 2. The framework consists of two main components: “Offline Sensor Selection” and “Online Diagnosis and Prognostics”.
I. Offline Sensor Selection: This stage occurs shortly after the habitat is deployed, during an initial period where degradation signals are collected with limited ground or crew assistance. The data collected during this stage forms a historical dataset, which is used to initialize the prognostic models before autonomous operation begins. We assume that the dataset contains multiple failure modes, although the failure modes are not labeled. The goal of this step is to (1) label the degradation signals into one of the failure modes, and (2) identify a subset of informative sensors for each failure mode. This step consists of two parts discussed below:
@@
    Feature Extraction: The onboard sensors generate high-dimensional, time-varying signals. To extract informative features for prognostics, we use a covariate-adjusted functional principal component analysis (CA-FPCA) method. CA-FPCA models the variation in each sensor signal while accounting for external covariates. In our setting, these covariates represent the underlying failure modes, which are unknown apriori. To estimate them, we first perform a K-means clustering on the sensor signals and use the resulting cluster labels as initial covariates.
@@
    Optimization: After feature extraction, we obtain a reduced set of features for each sensor. These features are called CA-FPC scores. We use these scores as predictors in a mixture of Gaussian regression (MGR) model. The MGR model estimates the natural logarithm of the time-to-failure, $\ln{TTF}$, as a function of the CA-FPC scores. The model assumes that this relationship ($\ln{TTF}$ vs CA-FPC scores) depends on the underlying failure mode. We fit the MGR model using the EM algorithm, which results in: (1) optimal failure mode labels for each sample, and (2) an optimal subset of informative sensors for each failure mode.
II. Online Diagnosis and Prognostics: Once, the failure mode labeling and sensor selection are complete, the habitat enters into autonomous mode. This stage uses real-time sensor data to (1) identify the active failure mode and (2) predict the RUL:
  1. 1.
    Diagnosis: Prior to the diagnosis step, the active failure mode is unknown therefore, all selected sensors are utilized in this step. We apply multivariate functional principal component analysis (MFPCA) to signals from all selected sensors to extract a compact set of features, called MFPC scores, that represent the joint behavior of all sensors. We classify the active failure mode by finding its K-nearest neighbors and assigning the most common failure mode among them.
  2. 2.
    Prognostics: Once the active failure mode is diagnosed, we recalculate the MFPC scores using only the informative sensors for that mode. A functional regression model uses these scores to predict the RUL.
4 Offline Sensor Selection